# Urn-driven random walks

### v2

@@
Urn-driven random walks
Srinivasan Balaji [^1] Department of Statistics, The George Washington University, Washington, D.C. 20052, U.S.A. Email: balaji@gwu.edu   Hosam Mahmoud [^2] Department of Statistics, The George Washington University, Washington, D.C. 20052, U.S.A. Email: hosam@gwu.edu
January 9, 2026
@@
keywords: Pólya urn, Friedman urn, random walk, exchangeability, Markov chain, Eulerian number, Catalan number.
Mathematical Subject Classification: 60G20, 50C05.
@@ 1 Introduction
If a state is not recurrent, it is called transient. Now, a random walk in $d$ dimensions arises by having a particle in $d$ dimensions changing one coordinate by a unit (positive or negative) at each point in discrete time, all moves being equally likely. For instance, in three dimensions, the particle chooses one of six possible displacements, which are namely, east, west, north, south, up or down, and each move has probability $\frac{1}{6}$. It is well known that the 1-dimensional and 2-dimensional random walks are recurrent, but the $d$-dimensional random walks, for $d\geq 3$, are all transient. This is a classic result first published by Pólya in 1920 (see [10]).
There is a variety of new ways to generate a random walk. For instance, the elephant random walk is in the vogue. This walk was introduced in 2004 in [11]. In recent years, numerous studies have been published on it; see for example [1, 2, 12].
@@ 2 Pólya random walk
Proof.
The joint distribution of the number of white and blue draws is the well-known Pólya distribution; see [8] or [9]. Pólya distribution is conveniently described in terms of Pochhammer’s symbol for the rising factorial, defined by
@@
where K:=(w+b1)!/((w1)!(b1)!)K:=(w+b-1)!/((w-1)!\,(b-1)!). This formula is a special case of Pólya distribution (also called the beta–binomial distribution), which first appeared in 1923 in [3]. Specialized to the case of $n$ white draws and $n$ blue draws, the formula follows from the exchangeability of the indicators of the draws. So, one can just lump all $n$ indicators of white draws at the beginning and count the number of sequences with $n$ white draws.
@@
We can count $Q_{2n}$ by outflanking any balanced expression of 2n22n-2 parentheses between an opening left and a closing right parenthesis. The first left parenthesis guarantees that, by the balancing of the outflanked expression, there will be more left than right parentheses, till position 2n12n-1 and the right parenthesis at position $2n$ balances the expression. The number of well balanced parentheses of length 2n22n-2 is well known to be the Catalan number ${2n-1\choose n-1}/n$; see [13, Section 1.3, pp. 12–14].
@@
note the doubling to take into account drawing sequences starting with $B$. Using the well-known asymptotics for the Catalan numbers [13, Section 2.3, pp. 35–37] and developments via the Stirling approximation of the factorials (as was done in the first part of the proof), we find
@@
Hence, state 0 in the Pólya random walk is null recurrent. ∎
@@ 3 Pólya random walk in higher dimensions
Figure 1 illustrates Remark 3.1. We can describe the path of Pólya walk in two dimensions in terms of ordered pairs $(x,y)\in\{W,B\}\times\{W,B\}$ specifying the colors drawn from the first and second urns, respectively.
@@
$\Big(\frac{1}{2}\times\frac{2}{3}\times\frac{1}{4}\times\frac{2}{5}\times\frac{3}{6}\times\frac{4}{7}\times\frac{5}{8}\times\frac{3}{0}\Big)\Big(\frac{2}{3}\times\frac{3}{4}\times\frac{4}{5}\times\frac{1}{6}\times\frac{5}{7}\times\frac{6}{8}\times\frac{7}{9}\times\frac{8}{10}\Big)$
@@ 4 Friedman random walk
In [4], the authors use an analytic method, to determine the probability distribution of $\widetilde{W}_{n}$, and for an urn starting with only one white ball find
@@
where $\big\langle{n\atop k}\big\rangle$ are the Eularian numbers of the first kind, that count the number of permutations of the set $\{1,2,\ldots,n\}$ with exactly $k$ ascents (see [7, Section 6.2]. For a broader discussion of the analytic method for treating urns see [5].
@@
The asymptotics of $\langle{n\atop k}\big\rangle$, for fixed $k$, as $n\to\infty$, had been known for a long time. More recently, Giladi and Keller [6] completed the spectrum by finding the asymptotics for all $k$. Actually, Giladi and Keller [6] develops the asymptotics of $A(n,k)$, which are the number of breaking points between the ascents. In other words, we have $A(n,k)=\big\langle{n\atop k+1}\big\rangle$. For our purpose, we are seeking the asymptotics of $\langle{2n\atop n+1}\big\rangle=A(2n,n)$. By the symmetry of the Eulerian numbers, we have A(2n,n)=A(2n,n1)A(2n,n)=A(2n,n-1). For even $n$, it is reported in [6] that
@@ X     $H_{r}\leftarrow W+B$
Author contributions statement
The two authors Srinivasan Balaji and Hosam Mahmoud have contributed equally to all the stages of the production of this manuscript: research, writing, and proofreading.
Funding declaration
This research is not funded.
Acknowledgment

### v3

@@
Abstract
@@
keywords:
Pólya urn, Friedman urn, random walk, exchangeability, Markov chain, Eulerian number, Catalan number
@@ 1 Introduction
If a state is not recurrent, it is called transient. Now, a random walk in $d$ dimensions arises by having a particle in $d$ dimensions changing one coordinate by a unit (positive or negative) at each point in discrete time, all moves being equally likely. For instance, in three dimensions, the particle chooses one of six possible displacements, which are namely, east, west, north, south, up or down, and each move has probability $\frac{1}{6}$. It is well known that the 1-dimensional and 2-dimensional random walks are recurrent, but the $d$-dimensional random walks, for $d\geq 3$, are all transient. This is a classic result first published by Pólya (see Pólya (1921)).
There is a variety of new ways to generate a random walk. For instance, the elephant random walk is in the vogue. This walk was introduced in Schütz and Trimper (2004). In recent years, numerous studies have been published on it; see for example Bercu and Laulin (2019); Bertoin (2022); Qin (2025).
@@ 2 Pólya random walk
Proof 2.2.
The joint distribution of the number of white and blue draws is the well-known Pólya distribution; see Johnson and Kotz (1977) or Mahmoud (2008). Pólya distribution is conveniently described in terms of Pochhammer’s symbol for the rising factorial, defined by
@@
where K:=(w+b1)!/((w1)!(b1)!)K:=(w+b-1)!/((w-1)!\,(b-1)!). This formula is a special case of Pólya distribution (also called the beta–binomial distribution), which first appeared in Eggenberger and Pólya (1923). Specialized to the case of $n$ white draws and $n$ blue draws, the formula follows from the exchangeability of the indicators of the draws. So, one can just lump all $n$ indicators of white draws at the beginning and count the number of sequences with $n$ white draws.
@@
We can count $Q_{2n}$ by outflanking any balanced expression of 2n22n-2 parentheses between an opening left and a closing right parenthesis. The first left parenthesis guarantees that, by the balancing of the outflanked expression, there will be more left than right parentheses, till position 2n12n-1 and the right parenthesis at position $2n$ balances the expression. The number of well balanced parentheses of length 2n22n-2 is well known to be the Catalan number ${2n-1\choose n-1}/n$; see (Stanley, 2015, Section 1.3, pp. 12–14).
@@
note the doubling to take into account drawing sequences starting with $B$. Using the well-known asymptotics for the Catalan numbers (Stanley, 2015, Section 2.3, pp. 35–37) and developments via the Stirling approximation of the factorials (as was done in the first part of the proof), we find
@@
Hence, state 0 in the Pólya random walk is null recurrent.
@@ 3 Pólya random walk in higher dimensions
Figure 1 illustrates Remark 2. We can describe the path of Pólya walk in two dimensions in terms of ordered pairs $(x,y)\in\{W,B\}\times\{W,B\}$ specifying the colors drawn from the first and second urns, respectively.
@@
$\Big(\frac{1}{2}\times\frac{2}{3}\times\frac{1}{4}\times\frac{2}{5}\times\frac{3}{6}\times\frac{4}{7}\times\frac{5}{8}\times\frac{3}{9}\Big)\Big(\frac{2}{3}\times\frac{3}{4}\times\frac{4}{5}\times\frac{1}{6}\times\frac{5}{7}\times\frac{6}{8}\times\frac{7}{9}\times\frac{8}{10}\Big)$
@@ 4 Friedman random walk
In Flajolet et al. (2006), the authors use an analytic method, to determine the probability distribution of $\widetilde{W}_{n}$, and for an urn starting with only one white ball find
@@
where $\big\langle{n\atop k}\big\rangle$ are the Eularian numbers of the first kind, that count the number of permutations of the set $\{1,2,\ldots,n\}$ with exactly $k$ ascents (see (Graham et al., 1994, Section 6.2)). For a broader discussion of the analytic method for treating urns see Flajolet et al. (2005).
@@
The asymptotics of $\langle{n\atop k}\big\rangle$, for fixed $k$, as $n\to\infty$, had been known for a long time. More recently, Giladi and Keller (1994) completed the spectrum by finding the asymptotics for all $k$. Actually, Giladi and Keller (1994) develops the asymptotics of $A(n,k)$, which are the number of breaking points between the ascents. In other words, we have $A(n,k)=\big\langle{n\atop k+1}\big\rangle$. For our purpose, we are seeking the asymptotics of $\langle{2n\atop n+1}\big\rangle=A(2n,n)$. By the symmetry of the Eulerian numbers, we have A(2n,n)=A(2n,n1)A(2n,n)=A(2n,n-1). For even $n$, it is reported in Giladi and Keller (1994) that
@@ X     $H_{r}\leftarrow W+B$
Acknowledgements.